# Enhancing Exchange Rate Forecasting with Explainable Deep Learning Models

### v1

@@ I Introduction
China’s growing role in the global supply chain, especially after its accession to the World Trade Organization (WTO), and the deepening economic ties between China and the United States, have made the RMB/USD exchange rate a focal point of global economic stability. Debates over the valuation of the RMB, particularly during periods of significant trade surpluses with the U.S., have led to multiple rounds of discussions on exchange rate policy. The RMB exchange rate reform in 2005 and the subsequent rounds of monetary policy adjustments by the Federal Reserve, especially during the 2007 financial crisis, further complicated the dynamics of the RMB/USD exchange rate. The “8.11 Exchange Rate Reform” in 2015, which introduced a more flexible exchange rate mechanism, and the intensified trade frictions since 2018, have put additional depreciation pressure on the RMB, making accurate forecasting of the RMB/USD exchange rate increasingly important [1, 2, 3, 4, 5].
@@
In recent years, there has been a notable shift towards using big data approaches in forecasting models, bypassing the need for complex mathematical modeling and allowing for more flexible model forms without predefined structures. The inherent complexity and non-linearity of exchange rate data have led to applying non-linear methods, such as chaos theory, non-parametric methods, and machine learning techniques, which have shown potential to improve forecasting accuracy. Studies like those of LeBaron and others have demonstrated that methods such as kernel ridge regression can significantly enhance the prediction of financial volatility, although some researchers, such as Mourer, have found that these methods do not always outperform simple autoregressive models in all contexts.
@@
Since the end of the Bretton Woods system, the transition to a floating exchange rate regime has posed significant challenges for managing risk, especially regarding the RMB/USD exchange rate. China’s integration into the global economy post-WTO accession, along with its deepening economic ties with the U.S., has made this exchange rate crucial for global economic stability. The 2015 reform of China’s exchange rate mechanism increased market-driven fluctuations, further intensified by trade tensions, which has heightened volatility. This evolving landscape has led to debates over the RMB’s valuation and spurred numerous policy discussions. Consequently, precise and reliable exchange rate forecasting has become essential for effective risk management in this volatile environment.
@@
Traditional exchange rate forecasting models, both theoretical and quantitative, have struggled with consistency and often neglect critical influencing factors. For example, theoretical models may fail to account for the real-time impact of policy changes and global economic shifts, while quantitative models may not fully capture the non-linear dynamics and intricate interactions of the variables involved. Recently, there has been a shift toward big data and machine learning approaches, which offer flexibility and improved accuracy in handling the complex, non-linear nature of exchange rate data [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. While some methods like kernel ridge regression have shown promise, their performance varies across different contexts, highlighting the ongoing challenges in exchange rate prediction.
@@
Machine learning models, particularly deep learning models, have increasingly been applied to predicting time series and economic variables [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. Despite their advantages in handling complex, non-linear data without requiring explicit assumptions about the underlying data distribution, these models are often criticized for their “black box” nature and lack of interpretability. Recent advancements, such as the application of Grad-CAM and attention mechanisms, have begun to address these issues, making it possible to visualize model predictions and understand the underlying decision-making processes. In related financial applications, researchers have combined advanced models like XGBoost with data balancing techniques including SMOTE to enhance predictive performance, as demonstrated by Chang et al. in their fraud detection study [60]. Their work not only showcased the effectiveness of this approach in fraud detection but also laid a foundation for developing robust models in various financial domains, including potentially exchange rate prediction [60]. However, applying these interpretability techniques has been mostly limited to fields like image recognition and natural language processing, with relatively few studies applying them to economic forecasting.
@@
Given the challenges of traditional models and the potential of machine learning approaches, this study seeks to explore the use of advanced deep learning models, including CNNs, RNNs, MLPs, and transformer-based architectures, for predicting the RMB/USD exchange rate. By incorporating a comprehensive set of features—drawn from economic indicators, trade data, and other currency pairs—and employing advanced feature selection techniques, this research aims to enhance predictive accuracy, identify the most relevant factors influencing exchange rate fluctuations, and enhance the interpretability of the model predictions.
@@ II-A Data Collection and Preprocessing
Macroeconomic Statistics: This includes key indicators like import/export values and short-term capital flows. While critical for decision-making, these statistics often suffer from delays, inaccuracies, and lack of predictive power. Branson’s (1975) asset portfolio approach to exchange rates, integrating purchasing power parity and risk-return models, is hampered by such data lags. Financial Market Trading Data: Includes real-time data like stock indices, concept stock indices, and exchange rates. These data are timely, transparent, and reflect market expectations, making them valuable for exchange rate predictions. Macroeconomic Variables: Covers indicators like M2 (monetary policy), spot interest rates, and price indices (inflation). Most data are daily; monthly data are treated as daily. This practice reflects real-world usage, where outdated data often guide predictions. Building upon recent advancements in financial data processing, we incorporated techniques inspired by Chang’s research [60] to significantly enhance our model’s processing efficiency, particularly in handling diverse data frequencies and potential redundancies.
We summarize the data we collected in Table. I. All data were sourced from the WIND Financial Terminal. Variables include the RMB/USD exchange rate and other relevant economic and financial indicators. Traditional regression models would struggle with multicollinearity and parameter estimation due to the extensive set of variables [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84]. However, machine learning models mitigate these issues [85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102]. Post-2015, the volatility of the USD/RMB exchange rate increased, leading us to focus on this period [103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 18, 114]. We applied z-score normalization to stabilize the model training process.
@@ II-C Models
We leverage the recently developed deep learning models, including TSMixer [115, 116, 117, 118, 119], FEDformer [120, 121, 122, 123, 124], LSTM [125, 126, 127], PatchTST [128, 129], TimesNet [130, 131, 132], Transformer [133, 134, 135], MLP [136, 137, 83, 138], TCN [139, 140, 141, 142], and iTransformer [143, 144, 145, 146, 147] for exchange rate prediction.
LSTM: We used three layers of LSTM with hidden dimensions of 32, 64, and 64. TCN: Applied in multi-variate time series prediction, using three layers of dilated temporal convolutions with residual connections. Kernel sizes were (3, 3, 5) with a dilation rate of (2, 2, 2). TSMixer [148, 149]: An MLP-based neural network with 5 layers, each capturing and mixing temporal features. Configured with a maximum feature dimension of 16 for balanced learning and efficiency. TimesNet: Combines convolutional layers to model local and global temporal patterns, transforming 1D time series into 2D tensors. The architecture includes 1 FFT block and 4 convolutional blocks. PatchTST [150]: Divides input sequences into patches, processed independently using 6 transformer layers with 8 heads per layer, and a patch size of 12. iTransformer: Trained with improved positional encoding and attention mechanisms, focusing on feature dimension attention. Consists of 4 transformer blocks, each with 2 layers and 8 attention heads. Fedformer: Configured with 3 encoder and 2 decoder layers, each using 8-head self-attention. The feature dimension is 256, with feedforward networks at 512. Transformer: Implemented with 6 layers in both encoder and decoder, using 8-head self-attention and a feature dimension of 512. Each layer has a feedforward network with an inner dimension of 2048, with positional encoding to capture temporal dependencies. MLP: Configured as a fully connected network with 3 hidden layers, each containing 128 neurons and ReLU activation for non-linearity.
@@ III-A Evaluation Metrics
In this study, model performance was evaluated using two key metrics: Mean Squared Error (MSE) and Mean Absolute Error (MAE). These metrics were selected to provide a thorough evaluation of the model’s predictive accuracy by quantifying the differences between the predicted and actual values. MSE emphasizes larger errors by squaring them, while MAE offers a straightforward measurement of the average error magnitude [151]. To ensure the reliability of the results, the metrics for each model were averaged across five-fold cross-validation, offering a more robust performance assessment.
@@ III-C Gradient Based Feature Importance Analysis
We applied grad-CAM [152, 153] to obtain visual explanations for decisions from our deep learning models, making them more transparent and explainable. grad-CAM generates heatmaps that capture the TSMixer’s focus during prediction on different features and time points, shown in Figure 2. The higher value (red) indicates the model is focusing more. We can clearly observe that the TSMixer model assigns significant weight to certain low-frequency macroeconomic variables, such as import and export volumes and trade amounts. However, the model appears to place even greater emphasis on trade-related data. It also shows a notable focus on fundamental data, such as the yield spread between U.S. 10-year Treasury bonds and the Federal Reserve’s interest rates, as well as the U.S. Consumer Price Index. The model assigns considerable importance to other currency pairs but does not prioritize stock indices like the CSI 300. This is visually evident in the figure, where these features stand out prominently. These findings suggest that variables with different frequencies play varying roles in exchange rate forecasting, but fundamental data remains crucial. This outcome also reinforces the reliability of using machine learning models for exchange rate prediction.

### v2

@@ I Introduction
China’s growing role in the global supply chain, especially after its accession to the World Trade Organization (WTO), and the deepening economic ties between China and the United States, have made the RMB/USD exchange rate a focal point of global economic stability. Debates over the valuation of the RMB, particularly during periods of significant trade surpluses with the U.S., have led to multiple rounds of discussions on exchange rate policy. The RMB exchange rate reform in 2005 and the subsequent rounds of monetary policy adjustments by the Federal Reserve, especially during the 2007 financial crisis, further complicated the dynamics of the RMB/USD exchange rate. The “8.11 Exchange Rate Reform” in 2015, which introduced a more flexible exchange rate mechanism, and the intensified trade frictions since 2018, have put additional depreciation pressure on the RMB, making accurate forecasting of the RMB/USD exchange rate increasingly important [1, 2, 3].
@@
In recent years, there has been a notable shift towards using big data approaches in forecasting models, bypassing the need for complex mathematical modeling and allowing for more flexible model forms without predefined structures. The inherent complexity and non-linearity of exchange rate data have led to the application of non-linear methods, such as chaos theory, non-parametric methods, and machine learning techniques, which have shown potential to improve forecasting accuracy. Studies like those of LeBaron and others have demonstrated that methods such as kernel ridge regression can significantly enhance the prediction of financial volatility, although some researchers, such as Mourer, have found that these methods do not always outperform simple autoregressive models in all contexts.
@@
Since the end of the Bretton Woods system, the transition to a floating exchange rate regime has posed significant challenges for managing risk, especially regarding the RMB/USD exchange rate. China’s integration into the global economy post-WTO accession, along with its deepening economic ties with the U.S., has made this exchange rate crucial for global economic stability. The 2015 reform of China’s exchange rate mechanism increased market-driven fluctuations, further intensified by trade tensions, which have heightened volatility. This evolving landscape has led to debates over the RMB’s valuation and spurred numerous policy discussions. Consequently, precise and reliable exchange rate forecasting has become essential for effective risk management in this volatile environment.
@@
Traditional exchange rate forecasting models, both theoretical and quantitative, have struggled with consistency and often neglect critical influencing factors. For example, theoretical models may fail to account for the real-time impact of policy changes and global economic shifts, while quantitative models may not fully capture the non-linear dynamics and intricate interactions of the variables involved. Recently, there has been a shift toward big data and deep learning approaches [4], which offer flexibility and improved accuracy in handling the complex, non-linear nature of exchange rate data. While some methods, like kernel ridge regression, have shown promise, their performance varies across different contexts, highlighting the ongoing challenges in exchange rate prediction.
@@
Machine learning models, particularly deep learning models, have increasingly been applied to predicting time series and economic variables. Despite their advantages in handling complex, non-linear data without requiring explicit assumptions about the underlying data distribution, these models are often criticized for their “black box” nature and lack of interpretability. Recent advancements, such as the application of Grad-CAM [5] and attention mechanisms, have begun to address these issues, making it possible to visualize model predictions and understand the underlying decision-making processes. However, applying these interpretability techniques has been mostly limited to fields like image recognition and natural language processing, with relatively few studies applying them to economic forecasting.
@@
Given the challenges of traditional models and the potential of machine learning approaches, this study seeks to explore the use of advanced deep learning models, including CNNs, RNNs, MLPs [6, 7, 8], and transformer-based [9] architectures, for predicting the RMB/USD exchange rate. By incorporating a comprehensive set of features—drawn from economic indicators, trade data, and other currency pairs—and employing advanced feature selection techniques, this research aims to enhance predictive accuracy, identify the most relevant factors influencing exchange rate fluctuations, and enhance the interpretability of the model predictions.
@@ II-A Data Collection and Preprocessing
Macroeconomic Statistics: This includes key indicators like import/export values and short-term capital flows. While critical for decision-making, these statistics often suffer from delays, inaccuracies, and lack of predictive power. Branson’s (1975) asset portfolio approach to exchange rates, integrating purchasing power parity and risk-return models, is hampered by such data lags. Financial Market Trading Data: Includes real-time data like stock indices, concept stock indices, and exchange rates. These data are timely, transparent, and reflect market expectations, making them valuable for exchange rate predictions. Macroeconomic Variables: Covers indicators like M2 (monetary policy), spot interest rates, and price indices (inflation). Most data are daily; monthly data are treated as daily. This practice reflects real-world usage, where outdated data often guide predictions. Building upon recent advancements in financial data processing, we incorporated techniques inspired by Chang’s research to significantly enhance our model’s processing efficiency, particularly in handling diverse data frequencies and potential redundancies.
We summarize the data we collected in Table. I. All data were sourced from the WIND Financial Terminal. Variables include the RMB/USD exchange rate and other relevant economic and financial indicators. Traditional regression models would struggle with multicollinearity and parameter estimation due to the extensive set of variables. However, machine learning models mitigate these issues. Post-2015, the volatility of the USD/RMB exchange rate increased, leading us to focus on this period. We applied z-score normalization to stabilize the model training process.
@@ II-C Models
We leverage the recently developed deep learning models, including TSMixer [10], FEDformer [11], LSTM [12], PatchTST [13], TimesNet [14], Transformer [9], MLP [15], TCN [16], and iTransformer [17] for exchange rate prediction.
LSTM: We used three layers of LSTM with hidden dimensions of 32, 64, and 64. TCN: Applied in multi-variate time series prediction, using three layers of dilated temporal convolutions with residual connections. Kernel sizes were (3, 3, 5) with a dilation rate of (2, 2, 2). TSMixer: An MLP-based neural network with 5 layers, each capturing and mixing temporal features. Configured with a maximum feature dimension of 16 for balanced learning and efficiency. TimesNet: Combines convolutional layers to model local and global temporal patterns, transforming 1D time series into 2D tensors. The architecture includes 1 FFT block and 4 convolutional blocks. PatchTST: Divides input sequences into patches, processed independently using 6 transformer layers with 8 heads per layer, and a patch size of 12. iTransformer: Trained with improved positional encoding and attention mechanisms, focusing on feature dimension attention. Consists of 4 transformer blocks, each with 2 layers and 8 attention heads. Fedformer: Configured with 3 encoder and 2 decoder layers, each using 8-head self-attention. The feature dimension is 256, with feedforward networks at 512. Transformer: Implemented with 6 layers in both encoder and decoder, using 8-head self-attention and a feature dimension of 512. Each layer has a feedforward network with an inner dimension of 2048, with positional encoding to capture temporal dependencies. MLP: Configured as a fully connected network with 3 hidden layers, each containing 128 neurons and ReLU activation for non-linearity.
@@ III-A Evaluation Metrics
In this study, model performance was evaluated using two key metrics: Mean Squared Error (MSE) and Mean Absolute Error (MAE). These metrics were selected to provide a thorough evaluation of the model’s predictive accuracy by quantifying the differences between the predicted and actual values. MSE emphasizes larger errors by squaring them, while MAE offers a straightforward measurement of the average error magnitude. To ensure the reliability of the results, the metrics for each model were averaged across five-fold cross-validation, offering a more robust performance assessment.
@@ III-C Gradient Based Feature Importance Analysis
We applied grad-CAM [5] to obtain visual explanations for decisions from our deep learning models, making them more transparent and explainable. grad-CAM generates heatmaps that capture the TSMixer’s focus during prediction on different features and time points, shown in Figure 2. The higher value (red) indicates the model is focusing more. We can clearly observe that the TSMixer model assigns significant weight to certain low-frequency macroeconomic variables, such as import and export volumes and trade amounts. However, the model appears to place even greater emphasis on trade-related data. It also shows a notable focus on fundamental data, such as the yield spread between U.S. 10-year Treasury bonds and the Federal Reserve’s interest rates, as well as the U.S. Consumer Price Index. The model assigns considerable importance to other currency pairs but does not prioritize stock indices like the CSI 300. This is visually evident in the figure, where these features stand out prominently. These findings suggest that variables with different frequencies play varying roles in exchange rate forecasting, but fundamental data remains crucial. This outcome also reinforces the reliability of using machine learning models for exchange rate prediction.